OpenStudy (anonymous):

I need some help really fast. Geometry, very important. I NEED HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! PLEASE!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

This is the question that I can't figure out.

OpenStudy (anonymous):

those exclamation marks tho

OpenStudy (anonymous):

I wasn't talking to anybody in particular, just trying to draw some attention.

OpenStudy (anonymous):

@amistre64 Can you help me?

OpenStudy (anonymous):

@monty5658 well that was helpful

OpenStudy (anonymous):

I know right, that x helped me a lot. NOT!

OpenStudy (anonymous):

I have no idea how to solve this but I'll try

OpenStudy (jdoe0001):

something tells me that's not the original graph....

OpenStudy (jdoe0001):

if I'm not mistaken, the bisecting line is really bisecting a 90 degree angle

OpenStudy (anonymous):

I swear that it is. I just have no idea on how to solve this.

OpenStudy (jdoe0001):

hmmm

OpenStudy (anonymous):

I just sort of feel like there is an infinite number of possibilities. The side of a triangle can be any length, and in this one we do not have enough information.

OpenStudy (anonymous):

Technically it is bisecting both the acute angle and a 90 degree angle, but I just need to know how to solve x from that informat

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

*information, my computer is acting really slow

OpenStudy (jdoe0001):

*information* yes, it's not really your machine, is the site itself

OpenStudy (anonymous):

I'm still so confused. This is my last problem, but I don't get it at all. Plus OpenStudy is being really slow for mew at least.

OpenStudy (jdoe0001):

notice what @monty5658 is suggesting

OpenStudy (anonymous):

*me

OpenStudy (anonymous):

I've never been taught what tan means. I am so bad at this. Is there anyway to break it down anymore I've never seen a problem where x is so haard to find.

OpenStudy (jdoe0001):

hmmm do you know the pythagorean theorem?

OpenStudy (anonymous):

Yes.

OpenStudy (jdoe0001):



OpenStudy (anonymous):

A little more sense. I'm going to review over everything and see if I can understand. Thanks for your help. I'll call you if I cafigure it out. Thanks for all of your help again.

OpenStudy (anonymous):

*can figure Sorry I have to close and open openstudy again because it is really slow for me. I can't even see what I've typed untI hit post.

OpenStudy (anonymous):

@monty5658, can you help me a bit more. I'm not good at solving long equations. I'm stuck at the point where at this mess: \[blank^{2}-blank^{2}=blank^{2}-blank^{2}\] Only did it that way to save time^

OpenStudy (anonymous):

@nikato , do you know how to solve the long equation that monty posted. I'm really confused.

OpenStudy (nikato):

Use foil to simplify the equation first

OpenStudy (anonymous):

I think that this is foil, I don't remember. First Outer Inner Last

OpenStudy (anonymous):

I don't know how to use it though.

OpenStudy (nikato):

But I'm not sure that's even right because you don't know if that's really a right angle. I think you would actually solve it using proportions And so Foil (10x-2)(10x-2) 10x(10x) + 10x(-2) + (-2)(10x) + (-2)(-2)

OpenStudy (anonymous):

What Nikato said..... I'm having computer issues. Nikato, my method only requires the heights to be the same ;)

OpenStudy (anonymous):

I'm still confused about your method @monty5658

OpenStudy (nikato):

Nvm, I'm sorry but Monty, your way does not work for this problem because it is not a right triangle. U have to solve with a proportion because all you know is an angle bisector http://www.cut-the-knot.org/triangle/AngleBisectorTheorem.shtml So 10x-2. 8x ------ =. ------ 7x. 5x+3

OpenStudy (anonymous):

I hate math so much. When will I ever use this is the future. In the future you need to be able to do basic math unless you are majoring in math. Nobody in their right mind that hates math as much as me would major in math, therefore I would never need this again in my life.

OpenStudy (nikato):

So cross multiply. And solve for x (10x-2)(5x+3) = 7x(8x)

OpenStudy (anonymous):

Agree with Nikato. My bad, I didn't see any angles on sheet and it appeared to be a set of right trianges.

OpenStudy (nikato):

Can you take it from here now?

OpenStudy (anonymous):

50x+30x, this is all I have, I don't know if this is even right.

OpenStudy (anonymous):

50x^2+30x, I don't know. I'm bad at this.

OpenStudy (nikato):

No, ,multiply everything out 50x^2 +20x -6=56x^2 -56x^2. -56x^2 --------------------------- -6x^2 +20x -6=0

OpenStudy (nikato):

It's a quadratic equation Factor a -2 -2(3x^2 -10x +3)=0

OpenStudy (nikato):

Factor -2(x-3)(3x-1)=0 So Solve x-3=0. Or 3x-1=0 2 solutions for x but only one is correct

OpenStudy (anonymous):

I think B, but I am not sure.

OpenStudy (nikato):

What's B?

OpenStudy (anonymous):

3x-1=0

OpenStudy (nikato):

So what did you get for x when you solved it?

OpenStudy (anonymous):

x=3?

OpenStudy (nikato):

Are you sure? 3x-1=0 +1. +1 ------_------ 3x=1 -- -- 3. 3 x=?

OpenStudy (anonymous):

.33333333333333333333333333...

OpenStudy (nikato):

I guess both solution work then. But won't it be easier with A because x=3 x-3=0 +3. +3 --------- x=3 And thanks for the testimonial

OpenStudy (anonymous):

Welcome. Thanks for your help.

OpenStudy (nikato):

No problem. So you actually have two solutions for this problem, 1/3 and 3

OpenStudy (nikato):

To check and make sure they are right, which already did, you can plug theses values into x to make sure they are right. If any side is negative, then that value of x is not right because you can't have a negative value for the length of a triangle